1.What is the concentration of HCl if it takes 63.0mL of 0.1 M NaOH to neutralize 140mL of the acid?


2. If it takes 58.5mL of 0.5 M potassium hydroxide to neutralize 135 mL of sulfuric acid, what is the concentration of the acid?


3. Describe the preparation of 2.00 L of 0.0500 M AgNO3 from the primary standard grade solid. Atomic masses: Ag = 108 N = 14 O =16



1
Expert's answer
2022-05-19T11:40:04-0400

1. What is the concentration of HCl if it takes 63.0 mL of 0.1 M NaOH to neutralize 140 mL of the acid?

Solution (1):

Balanced chemical equation:

HCl + NaOH → NaCl + H2O

According to stoichiometry:

Moles of HCl = Moles of NaOH

Molarity of HCl × Volume of HCl = Molarity of NaOH × Volume of NaOH

Therefore,

Molarity of HCl = (Molarity of NaOH × Volume of NaOH) / (Volume of HCl)

Molarity of HCl = (0.1 M × 63.0 mL) / (140 mL) = 0.045 M

Molarity of HCl = 0.045 M

Answer (1): The concentration of HCl is 0.045 M


2. If it takes 58.5 mL of 0.5 M potassium hydroxide to neutralize 135 mL of sulfuric acid, what is the concentration of the acid?

Solution (2):

Balanced chemical equation:

H2SO4 + 2KOH → K2SO4 + 2H2O

According to stoichiometry:

Moles of H2SO4 = Moles of KOH / 2

2 × Molarity of H2SO4 × Volume of H2SO4 = Molarity of KOH × Volume of KOH

Therefore,

Molarity of H2SO4 = (Molarity of KOH × Volume of KOH) / (2 × Volume of H2SO4)

Molarity of H2SO4 = (0.5 M × 58.5 mL) / (2 × 135 mL) = 0.108 M

Molarity of H2SO4 = 0.108 M

Answer (2): The concentration of H2SO4 is 0.108 M


3. Describe the preparation of 2.00 L of 0.0500 M AgNO3 from the primary standard grade solid. Atomic masses: Ag = 108 N = 14 O =16.

Solution (3):

Molarity = Moles of solute / Liters of solution

Therefore,

Moles of AgNO3 = Molarity × Liters of solution

Moles of AgNO3 = (0.0500 M) × (2.00 L) = 0.1 mol


Molar mass of AgNO3 = Ar(Ag) + Ar(N) + 3×Ar(O) = 108 + 14 + 3×16 = 170

Molar mass of AgNO3 = 170 g/mol


Moles = Mass / Molar mass

Therefore,

Mass of AgNO3 = Moles of AgNO3 × Molar mass of AgNO3

Mass of AgNO3 = (0.1 mol) × (170 g/mol) = 17 g

Mass of AgNO3 = 17 g


Thus, to prepare such a solution, you need to take 17 g of silver nitrate (AgNO3), dissolve it in water and dilute to 2.00 liters